4. Compute the value of 21410-14610 in binary and write your final answer in octal
1
Expert's answer
2020-06-01T10:36:14-0400

To do binary subtraction between two given numbers, first, we have to convert the given two numbers into binary. Since the given two numbers are not in binary so we will convert these two numbers into binary. The binary representation of the two given numbers as follows as:







Now:

We will do subtract between these two binary numbers. To do subtract between two binary numbers, we will follow the following rules:

1 - 0 = 1

0 - 0 = 0

1 - 1 = 0

0 - 1 = 1 with a borrow of 1

Using these rules, we subtract the two binary numbers 101001110100010 and 11100100010010 as follows as:




Now:

To convert the binary subtraction result into an octal number, we divide the binary number into three bits group from the right side of the binary number and then convert each three bits group into decimal numbers. The final answer is in the form of an octal number. The octal representation of the subtraction result is as follows as: